Page 2 II. (7) Use the Laplace transform to solve the IVP y - 5y + 6y = 8(t-1), y(0) = 0,0) = 0, where the right hand side

f(t)=L-{F(s) Table of Laplace Transforms F(s)=L{()} f(t)= L-{F(s) F(s)=L{f(t)} 1. 2. et s-a 3. r, n=1,2,3,... P+1 5. Vi a
Answer #1

IYP is tet Then we get y => 5 Given 7- 5y+by = $(t-1), 4(0) =o=y(o). taplace transform of y(t) is 4 {y(t)} F(s) taking Lap